OpenStudy (vcabral1134):

Help me plz will fan and medal! Write a recursive rule for the sequence. a^n = −9+ 2n a. a^1 = −9,a^n = a^n − 1 − 9 b. a^1 = −9,a^n = a^n − 1 + 2 c. a^1 = −7,a^n = a^n − 1 − 9 d. a^1 = −7,a^n = a^n − 1 + 2

OpenStudy (vcabral1134):

I eliminated C and D.

OpenStudy (saurabh135):

d is the correct answer

OpenStudy (vcabral1134):

How?? @Saurabh135

OpenStudy (saurabh135):

see if we put n=1 in above equation, then a^1 will be equal to -9+2*1 = -7

OpenStudy (saurabh135):

and if we put n-1 instead of n in above equation then a^n-1 = -9 + 2*(n-1) = -9 + 2n -2 and subtracting these equations we get a^n - a^n-1 = 2 which implies a^n = a^n-1 + 2